Need help with this word problem if some one would have the time please, and thank you!!
3. You recently started the paperwork to purchase your new home, and you were just notified that you can move into the house in 2 weeks. You decide to hire a moving company, but are unsure which company to choose. You search online and are interested in contacting two companies, Heavy Lifters and Quick Move, to discuss their rates. Heavy Lifting charges an $80 fee plus $35 per hour. Quick Move charges $55 per hour with no additional fees.
a) Which mover provides a better deal for 2 hours of work? How did you arrive at your answer?
b) Which mover provides a better deal for 15 hours of work? How did you arrive at your answer?
c) For what values h (hours) does Quick Move offer the better deal? Express your answer as an inequality. Explain how you reached your answer.


Answer by stanbon(75887)

Heavy Lifting charges an $80 fee plus $35 per hour.
CostA = 80 + 35x
---------------------
Quick Move charges $55 per hour with no additional fees.
CostB = 55x
----------------------
a) Which mover provides a better deal for 2 hours of work? How did you arrive at your answer?
A(2) = 80+35*2 = 150
B(2) = 55*2 = 110
Ans: B
---------------------
b) Which mover provides a better deal for 15 hours of work? How did you arrive at your answer?
A(15) = 80 + 35*15 = 605
B(15) = 55*15 = 825
Ans: A
--------------------
c) For what values h (hours) does Quick Move(B) offer the better deal? Express your answer as an inequality. Explain how you reached your answer.
INEQUALITY:
55x < 80 + 35x
20x < 80
x < 4 hours
===================
Cheers,
Stan H.